OpenStudy (anonymous):

without graphing, determine the equation whose graph intersects the graph of -6x + 3y = 11 exactly once. how do i get my answer?

OpenStudy (unklerhaukus):

the only way to touch a parabola a a single point is to touch it at the vertex

OpenStudy (anonymous):

yes, but how do I figure this out without graphing? my choices are: i) -6x + 3y = 13 ii) -24x + 12y = 44 iii) -4x + 3y = 11

OpenStudy (anonymous):

also iv) none, lol.

OpenStudy (jagatuba):

iii

OpenStudy (jagatuba):

y-intercept is 11/3

OpenStudy (anonymous):

how did you arrive there please.=

OpenStudy (unklerhaukus):

look at the right hand side of equations

OpenStudy (unklerhaukus):

the number is the same

OpenStudy (anonymous):

I must be looking at this wrong.. is (-6x + 3y = 11) y=6/3+11 in y=mx+b format?

OpenStudy (jagatuba):

{y=2x+11/3, y=(4x)/3+11/3}

OpenStudy (jagatuba):

Y-intercept on both equations is 11/3 and the slope is different therefore the lines intersect at (0, 1/3)

OpenStudy (jagatuba):

(0, 11/3) sorry

OpenStudy (anonymous):

{y=2x+11/3, y=(4x)/3+11/3} ; why is the mx = 2x and 4x???????

OpenStudy (jagatuba):

2x+11/3=(4x)/3+11/3